OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (4, -1) and perpendicular to the line whose equation is 2x - y - 7 = 0

OpenStudy (anonymous):

Line perpendicular to the given line will have negative reciprocal slope from the given line that it y = 2x -7 m = 2 slope of perpendicular line = -1/2 perpendicular line passes through (4,-1) so using point slope equation form y - y1 = m(x-x1) y -(-1) = (-1/2)(x -4) y +1 = -x/2 +2 y = -x/2 +1

OpenStudy (anonymous):

The answers are y = -2x y = -1/2x y =1/2x y = 2x Which one is equal to y = -x/2 +1

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

y = -1/2 x

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Anytime :)